A sample of helium occupies 535 ml at 988 torr and 25°C. If the sample is transferred to a 1.05 L flask at 25°C, what will be the gas pressure in the flask?


1
Expert's answer
2021-06-07T03:11:40-0400

We know that ideal gas equation is PV = nRT


Where;

T = Temperature = 25.5 oC = 25.5+273 = 298.5 K

P = pressure = 988 mm Hg = 988/760 atm = 1.3 atm


n = No . of moles = ?


R = gas constant = 0.0821 L atm / mol - K


V= Volume of the gas = 535 mL = 0.535 L


Plug the values we get n = (PV) / (RT)


= 0.0284 moles


So the number of moles of He gas is 0.0284 moles


Now we have to calculate the pressure when it is transferred to 1.05 L


We know that ideal gas equation is PV = nRT


Where


T = Temperature = 25.5 oC = 25.5+273 = 298.5 K


P = pressure = ?


n = No . of moles = 0.0284 mol


R = gas constant = 0.0821 L atm / mol - K


V= Volume of the gas = 1.05 L


Plug the values we get P = (nRT) / V


= ( 0.0284x0.0821x298.5) / 1.05


= 0.662 atm


= 0.662x760 mm Hg

since 1atm = 760 mm Hg

= 503.4 mm Hg